Your weekly paycheck is 25 percent less than your coworker's. Your two paychecks total 660. Find the amount of each paycheck.
Found 2 solutions by Maths68, musef:
Answer by Maths68(1474)

Let
My paycheck amount = a
25% of my paycheck amount = a * 25%
CoWorker's paycheck amount = a + (a * 25%)=a+(25a/100)=a+(a/4)=(4a+a)/4=5a/4
Two paychecks = 660
My paycheck amount + CoWorker's paycheck amount = 660
a+5a/4=660
Multiply by 4 both sides of the equation
4a+5a=2640
9a=2640
a=2648/9
a=293.33
my paycheck amount = 293.33
my paycheck amount + CoWorker's paycheck amount = 660
293.33 + CoWorker's paycheck amount = 660
CoWorker's paycheck amount = 660-293.33
CoWorker's paycheck amount = 366.67

Answer by musef(1)


the above solution is WRONG!, pls.

Consider your co-worker's paycheck as = a
your paycheck = 0.75a

now solve!